It is found experimentally that 13.6 eV energy is required to separate a hydrogen atom into a proton and an electron. Compute the orbital radius and the velocity of the electron in a hydrogen atom.

Grade:11

2 Answers

opakunleabdulrahman
6 years ago
R=e^2. (1.6×10^19)^2 ------- = --------------------- 8πE0E 8π[8.85×10^—12)(–2.2×10^–18). ans 5.3×10^-18(2). e. 1.6×10^–18 --------------- = ------------------ ans2.2×10^6 √4πE0mr. 4π×8.86×10^12×9.1×10^31×5.3×10^–11
Aashutosh singh
6 years ago
Total energy in H atom = 13.6 eV= -2.2*10^-18 Total energy of an atom = e^2/8π£°rBy above equationse^2/8π£°r= -2.2 * 10^-18Therefore r = - e^2/8π£°E On solving r = 5.3*10^-11